OpenStudy (richyw):

Classifying singular points.\[f(z)=\frac{1}{z(z-1)}\]

OpenStudy (richyw):

ok so my singular points are z = 0 and z = 1. I am trying to write the Laurent series of f(z) about z=0 in the region 0<|z|<1

OpenStudy (richyw):

nvm. I figured out my problem.

OpenStudy (richyw):

I forgot that \[\frac{1}{1-z}=\sum_{n=0}^\infty z^n\]for some reason thought it was 1/(z-1)